### v1

# Learning Degenerate Manifolds of Frustrated Magnets with Boltzmann Machines

Abstract
We show that Restricted Boltzmann Machines (RBMs) provide a flexible generative framework for modeling spin configurations in disordered yet strongly correlated phases of frustrated magnets. As a benchmark, we first demonstrate that an RBM can learn the zero-temperature ground-state manifold of the one-dimensional ANNNI model at its multiphase point, accurately reproducing its characteristic oscillatory and exponentially decaying correlations. We then apply RBMs to kagome spin ice and show that they successfully learn the local ice rules and short-range correlations of the extensively degenerate ice-I manifold. Correlation functions computed from RBM-generated configurations closely match those from direct Monte Carlo simulations. For the partially ordered ice-II phase—featuring long-range charge order and broken time-reversal symmetry—accurate modeling requires RBMs with uniform-sign bias fields, mirroring the underlying symmetry breaking. These results highlight the utility of RBMs as generative models for learning constrained and highly frustrated magnetic states.
I introduction
Machine learning (ML) has seen rapid adoption across physics and chemistry in recent years [1, 2, 3, 4, 5, 6]. Supervised approaches now play a central role in predicting materials properties, constructing phase diagrams, and classifying quantum states. A notable success is the development of ML-based interatomic potentials, which now underpin ab initio molecular dynamics (MD) simulations [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] and a growing range of condensed-matter applications [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. In parallel, unsupervised learning has become an indispensable tool for uncovering structure in unlabeled data: methods ranging from PCA to modern generative models such as variational autoencoders (VAEs) extract dominant modes, reveal latent variables, and capture complex many-body correlations. Together, these ML techniques have enabled the identification of new phases of matter, the detection of emergent features in quantum systems, and the discovery of hidden organizing principles in large-scale simulation and experimental datasets.
A prototypical unsupervised generative model is the Restricted Boltzmann Machine (RBM) [34, 35, 36, 37]. An RBM is designed to learn the probability distribution underlying a dataset by reconstructing inputs from a set of latent (hidden) variables. Its architecture consists of a visible layer representing the data and a hidden layer encoding latent features, with full connectivity between layers but no intra-layer couplings—hence the “restricted” bipartite structure. This design enables efficient evaluation of visible–hidden correlations and supports stochastic training procedures, such as contrastive divergence, that approximate maximum-likelihood learning. By capturing a compact probabilistic representation of the data, RBMs provide powerful tools for generative modeling, feature extraction, and dimensionality reduction.
As generative models, RBMs have shown a remarkable ability to reproduce equilibrium properties of many-body systems [38, 39, 40, 41, 42]. Trained on Monte Carlo configurations of the square-lattice Ising model, for instance, an RBM can learn the underlying equilibrium distribution and generate new samples that accurately reproduce heat-capacity, magnetization, and susceptibility. Their generative power has also inspired hybrid Monte Carlo schemes in which RBMs propose nonlocal, cluster-type updates that improve sampling efficiency [43, 44, 45, 46, 47]. Beyond classical systems, RBMs have also been employed as expressive variational ansätze for quantum many-body states [48, 49, 50, 51]. In this context, the RBM represents the complex wave function amplitudes in a compact latent space, enabling efficient variational optimization of ground states. These successes reflect a deep structural analogy between RBMs and equilibrium statistical mechanics: the RBM energy function closely mirrors a classical spin Hamiltonian. Consequently, RBMs serve both as practical computational tools and as conceptual bridges linking modern machine learning with statistical and quantum many-body physics.
In this work, we apply RBMs to a qualitatively different class of classical statistical-physics problems. Instead of focusing on ordered phases with relatively simple spatial correlations, we investigate whether an RBM can learn the highly correlated yet intrinsically disordered spin-liquid phases that arise in strongly frustrated magnets. In such systems, pairwise spin interactions cannot be simultaneously satisfied—whether due to special lattice geometries such as corner-sharing triangles or tetrahedra, or due to competing exchange interactions. This frustration leads to an extensively degenerate ground-state manifold, characterized by nontrivial local constraints and intricate correlation patterns that persist without conventional symmetry breaking.
Our central goal is to determine whether an RBM can faithfully approximate the underlying probability distribution of these spin-liquid ensembles, despite the absence of long-range order and the subtle constraints governing their low-energy manifold. More broadly, we seek to assess the capabilities and limitations of RBMs as generative models for frustrated, highly degenerate classical systems — examining whether their compact latent representations can capture emergent constraints, gauge-like degrees of freedom, and long-range correlations characteristic of classical spin liquids.
As concrete case studies, we apply the RBM framework to two paradigmatic examples: the extensively degenerate ground-state manifold of the one-dimensional axial next-nearest-neighbor Ising (ANNNI) model [52, 53, 54, 55], and the correlated spin configurations of two-dimensional kagome spin-ice systems [56, 57, 58, 59]. These systems provide complementary testbeds—one with one-dimensional, constraint-driven degeneracy, and the other with two-dimensional, Coulomb-phase-like correlations—allowing us to probe the robustness of RBM-based generative modeling across distinct forms of frustration.
The remainder of this paper is organized as follows. In Sec. II, we briefly review the RBM architecture and training procedures relevant to probabilistic modeling of classical spin configurations. Section III presents our results for the ANNNI model, highlighting the RBM’s ability to learn constraint-dominated ensembles. Section IV then examines the kagome spin-ice system, focusing on the RBM’s capacity to reproduce algebraic correlations and divergence-free conditions. We conclude in Sec. V with a summary of our findings and a discussion of broader implications for ML-based generative modeling of frustrated and highly degenerate statistical systems.
II Restricted Boltzmann Machines
Restricted Boltzmann machines (RBMs) are energy-based generative models defined on a bipartite graph comprising visible and hidden units. Introduced as a tractable specialization of the fully connected Boltzmann machine, RBMs eliminate all intra-layer couplings, thereby yielding factorized conditional distributions and enabling efficient Gibbs sampling and approximate maximum-likelihood training. This architectural simplification makes RBMs well suited for modeling high-dimensional probability distributions and has contributed to their widespread use in machine-learning applications since the mid-2000s.
Figure 1: Schematic diagram of a restricted Boltzmann machine (RBM) in an Ising-spin representation. The visible spins $\bm{\sigma}$ (green) are coupled to the hidden spins $\bm{\tau}$ (yellow) through the weight matrix $\bm{W}$. The local fields $\mathbf{b}$ and $\mathbf{b}^{\prime}$ appearing in the RBM Hamiltonian are indicated by external boxes connected to the visible and hidden layers, respectively.
From a physics perspective, an RBM may be viewed as a bipartite spin-glass model defined on two sets of Ising variables, $\bm{\sigma}=(\sigma_{1},\sigma_{2},\cdots,\sigma_{N})$ and $\bm{\tau}=(\tau_{1},\tau_{2},\cdots,\tau_{M})$; see Fig. 1. Here $N$, $M$ are the number of spins (neurons) in the visible and hidden layers, respectively. Every visible spin couples to every hidden spin through the edges of the bipartite graph. The associated coupling strengths are collected in a weight matrix $\mathbf{W}$ with vanishing diagonal, where $W_{ij}$ parameterizes the interaction between $\sigma_{i}$ and $\tau_{j}$. Two sets of local fields, $\bm{b}=(b_{1},b_{2},\cdots,b_{N})$ and $\bm{b}^{\prime}=(b^{\prime}_{1},b^{\prime}_{2},\cdots,b^{\prime}_{M})$, act on the visible and hidden layers, respectively. The joint distribution realized by the RBM is an “equilibrium” Boltzmann distribution,
$\pi_{\bm{\theta}}(\bm{\sigma},\bm{\tau})=\frac{1}{\mathcal{Z}_{\bm{\theta}}}e^{-E_{\bm{\theta}}(\bm{\sigma},\bm{\tau})},$ (1)
where $\bm{\theta}=\{W_{ij},b_{i},b^{\prime}_{j}\}$ denotes the full set of parameters, $\mathcal{Z}_{\bm{\theta}}$ is the partition function, and the effective energy takes the form
$E_{\bm{\theta}}(\bm{\sigma},\bm{\tau})=-\sum_{i=1}^{N}\sum_{j=1}^{M}W_{ij}\sigma_{i}\tau_{j}-\sum_{i=1}^{N}b_{i}\,\sigma_{i}-\sum_{j=1}^{M}b^{\prime}_{j}\,\tau_{j}.$
The bipartite structure also enables efficient nonlocal dynamical updates. Because the state of each node is sampled from a nonlinear function of its inputs (its “activation”), and because nodes within the same layer are conditionally independent, one may update all visible units simultaneously given the hidden layer, and vice versa. This layer-wise sampling can be implemented using fast linear-algebra operations. In particular, for any fixed parameter set $\bm{\theta}$, the conditional distributions $p_{\bm{\theta}}(\bm{\sigma}\,|\,\bm{\tau})$ and $p_{\bm{\theta}}(\bm{\tau}\,|\,\bm{\sigma})$ factorize over sites and are obtained directly via Bayes’ rule:
$p_{\bm{\theta}}(\bm{\sigma}\big|\bm{\tau})=\prod_{i=1}^{N}p_{\bm{\theta}}(\sigma_{i}\big|\bm{\tau}),\quad p_{\bm{\theta}}(\bm{\tau}\big|\bm{\sigma})=\prod_{j=1}^{M}p_{\bm{\theta}}(\tau_{j}\big|\bm{\sigma})$ (3)
with the single-site probabilities
$p_{\bm{\theta}}(\sigma_{i}=1\big|\bm{\tau})={\rm sigm}\Bigl(\sum_{j}W_{ij}\tau_{j}+b_{i}\Bigr),$
$p_{\bm{\theta}}(\tau_{j}=1\big|\bm{\tau})={\rm sigm}\Bigl(\sum_{i}W_{ij}\sigma_{i}+b^{\prime}_{j}\Bigr)$ (4)
where ${\rm sigm}(x)=(1+e^{-x})^{-1}$ denotes the logistic (sigmoid) activation function.
As a generative model, the RBM represents a target probability distribution through the marginal obtained by integrating out the hidden spins:
$\pi_{\bm{\theta}}(\bm{\sigma})=\sum_{\bm{\tau}}\pi_{\bm{\theta}}(\bm{\sigma},\bm{\tau})=\frac{1}{\mathcal{Z}_{\bm{\theta}}}\sum_{\bm{\tau}}e^{-E_{\bm{\theta}}(\bm{\sigma},\bm{\tau})}.$ (5)
Integrating out the hidden layer induces nontrivial, effectively higher-order interactions among the visible spins. Owing to the bipartite structure of the RBM, this marginalization can be performed analytically, yielding
$\pi_{\bm{\theta}}(\bm{\sigma})=\frac{1}{\mathcal{Z}_{\bm{\theta}}}e^{-E_{\bm{\theta}}(\bm{\sigma})},$ (6)
with an effective free energy of the visible spins:
$E_{\bm{\theta}}(\bm{\sigma})=-\sum_{i}b_{i}\sigma_{i}-\sum_{j}\ln\left(1+e^{b^{\prime}_{j}+\sum_{i}W_{ji}\sigma_{i}}\right)$ (7)
The second term encapsulates the influence of the hidden spins and generates highly nonlinear couplings among the visible units, enabling the RBM to model complex correlated manifolds. Training the RBM amounts to adjusting $\bm{\theta}$ so that $\pi_{\bm{\theta}}(\bm{\sigma})$ approximates a desired target distribution. Once trained, new samples are produced by alternating Gibbs updates of the visible and hidden layers, an efficient sampling strategy afforded by the bipartite architecture.
In the present context, the RBM is tasked with approximating the Boltzmann distribution of a classical spin Hamiltonian $\mathcal{H}_{\rm phys}(\bm{\sigma})$,
$\pi_{\rm phys}(\bm{\sigma})=\frac{e^{-\beta\mathcal{H}_{\rm phys}(\bm{\sigma})}}{\mathcal{Z}_{\rm phys}},$ (8)
where $\beta=1/k_{B}T$ and $\mathcal{Z}_{\rm phys}=\sum_{\bm{\sigma}}e^{-\beta\mathcal{H}_{\rm phys}(\bm{\sigma})}$. The training objective is to minimize the Kullback–Leibler divergence between $\pi_{\bm{\theta}}$ and $\pi_{\rm phys}$. The parameters evolve according to the stochastic gradient update
$\bm{\theta}^{(n+1)}=\bm{\theta}^{(n)}+\eta\left[\left\langle-\frac{\partial E_{\bm{\theta}}}{\partial\bm{\theta}}\right\rangle_{\rm data}-\left\langle-\frac{\partial E_{\bm{\theta}}}{\partial\bm{\theta}}\right\rangle_{\rm model}\right],\quad$ (9)
[…]
III ANNNI model
The anisotropic next-nearest-neighbor Ising (ANNNI) model [52, 53, 54, 55] is a canonical classical spin system for studying frustrated magnetism, modulated phases, and commensurate-incommensurate transitions. By supplementing the ferromagnetic Ising model with an antiferromagnetic next-nearest-neighbor interaction along a single spatial axis, the model introduces competition between couplings of different signs and ranges. This interaction-induced frustration—present even on bipartite lattices—stabilizes spatially modulated configurations that do not arise in the nearest-neighbor Ising model. Because the underlying symmetry remains purely Ising-like and geometric frustration is absent, the ANNNI model provides a clean theoretical setting for analyzing complex ordering phenomena generated solely by competing interactions.
Figure 2: (a) Competing interactions in the 1D ANNNI chain: ferromagnetic nearest-neighbor bonds $J_{1}$ and antiferromagnetic next-nearest-neighbor bonds $J_{2}$. (b) Ferromagnetic ground state for $\kappa=J_{2}/J_{1}<1/2$, and (c) the anti-phase $\langle 2,2\rangle$ state for $\kappa>1/2$.
In higher dimensions, such competition produces a finite-temperature phase diagram containing an infinite hierarchy of commensurate modulated phases. Between the uniform ferromagnet and the period-2 antiphase, the model supports a dense sequence of ordered structures of the form $\uparrow^{m}\downarrow^{n}$, each characterized by a rational modulation wave vector. These states form a devil’s staircase: a fractal hierarchy of locked plateaus that proliferate near the $T=0$ boundary separating the uniform and period-2 phases, and expand into a characteristic fan-shaped region of commensurate order at finite temperature. At still higher temperatures, this staircase gives way to an incommensurate (floating) phase in which the modulation wave vector varies continuously.
As a warm-up example for the RBM application, we focus on the one-dimensional ANNNI chain with Hamiltonian
$\mathcal{H}_{\rm ANNNI}(\bm{\sigma})=-J_{1}\sum_{i=1}^{N}\sigma_{i}\sigma_{i+1}+J_{2}\sum_{i=1}^{N}\sigma_{i}\sigma_{i+2},$ (10)
where $J_{1}>0$ is the ferromagnetic nearest-neighbor interaction and $J_{2}>0$ is the antiferromagnetic next-nearest-neighbor coupling along the axial direction; see Fig. 2(a) for a schematic of the model. Periodic boundary condition is assumed. Although the 1D model lacks true long-range order at any finite temperature, its zero-temperature phase diagram already captures the essential effects of frustration. For $\kappa=J_{2}/J_{1}<1/2$, the ground state is the uniform ferromagnetic configuration $\uparrow\uparrow\uparrow\cdots$, determined by the dominant nearest-neighbor term. For $\kappa>1/2$, frustration stabilizes a modulated period-2 antiphase of the form $\uparrow\uparrow\downarrow\downarrow\cdots$. At the special point $\kappa=1/2$, the domain-wall energy vanishes and the model develops a macroscopically large manifold of degenerate ground states: any spin sequence composed solely of domains of length one or two has identical energy density, reflecting the collapse of the effective tension between adjacent spin blocks.
Here we are interested in the special point $\kappa=1/2$, where the competition between $J_{1}$ and $J_{2}$ leads to extensive degeneracy. Although the $\kappa=1/2$ at $T=0$ is not a critical point in the sense of a divergent correlation length, it nevertheless exhibits highly nontrivial short-range correlations rooted in the macroscopic degeneracy. At $T=0$ and $\kappa=1/2$, any spin configuration with no three consecutive spins of the same sign is a ground state. The number of such configurations grows as $\mu^{N}$ [60], where $\mu=(1+\sqrt{5})/2$ is the golden ratio, so the residual entropy per spin is $S_{0}=\ln\mu$.
In the natural ensemble in which all ground states are weighted equally, the longitudinal spin-spin correlator is short-ranged and oscillatory [61]. A transfer-matrix analysis over the constrained ground-state manifold shows that, in the thermodynamic limit, $C(r)=A\cos(2\pi r/3+\phi)e^{-r/\xi}$ with a finite correlation length $\xi=1/\ln(\frac{3+\sqrt{5}}{2})\approx 1.04$, and the amplitudes/phases $A$ and $\phi$ that differ between even and odd sites.
Figure 3: Plot of $\varepsilon_{W}$ during training of the ice-II phase RBM. Early oscillatory behavior indicates rapid adjustment of model parameters, while decay of $\varepsilon_{W}$ at later iterations indicates convergence of the gradient descent procedure.
Figure 4: Comparison of the spin-spin correlation function $C(r)$ obtained from Markov-chain Monte Carlo sampling and from RBM-generated configurations for the ANNNI chain at the multiphase point $\kappa=1/2$ and $T=0$.
Since we aim to train an RBM at the multiphase point $\kappa=1/2$ and $T=0$, the corresponding target distribution is simply uniform over the degenerate ground-state manifold. More explicitly, $\pi_{\rm ANNNI}(\bm{\sigma};\,\kappa=1/2,T=0)=1/\mathcal{Z}$ for any spin configuration containing no three consecutive spins of the same sign, where $\mathcal{Z}\sim\mu^{N}$ is simply the number of such degenerate states. To generate samples from this ensemble, we use standard Markov-chain Monte Carlo (MC) simulations of the ANNNI chain at zero temperature with single-spin Metropolis updates. Because all domain walls separating the ferromagnetic and period-2 antiphase segments have zero energy cost at $\kappa=1/2$, the acceptance rate for these updates remains high even at $T=0$, enabling efficient exploration of the degenerate manifold.
Figure 5: (a) Ice-rule configurations (2-in/1-out and 1-in/2-out) for three spins on a triangle unit of the kagome lattice, each carrying an emergent magnetic charge of $Q=\pm 1$ unit. (b) Elementary excitations consist of defect triangles with 3-in or 3-out spin configurations, corresponding to magnetic charges of $Q=\pm 3$. (c) Example of an ice-I state. Small red and blue circles mark triangles that satisfy the ice-rule constraint and carry magnetic charges $Q=+1$ and 1-1, respectively, while large red and blue circles denote defect triangles with $Q=\pm 3$ that violate the constraint. The inset illustrates a single triangular unit cell with the three easy-axis directions $\hat{\mathbf{e}}_{1,2,3}$ of the kagome lattice. (d) Example of the charge-ordered ice-II phase, in which the $\pm 1$ magnetic charges form a staggered pattern that breaks the $Z_{2}$ symmetry relating up- and down-triangles.
Spin configurations $\{\tilde{\bm{\sigma}}^{(s)}\}$, $s=1,2,\cdots,N_{\rm data}$, generated from the MC simulations are then used to train the RBM via the stochastic gradient descent update in Eq. (9). In particular, the data-dependent gradient term $\langle\partial E_{\bm{\theta}}/\partial\bm{\theta}\rangle_{\rm data}$ in evaluated by clamping the visible spins to the MC-sampled spin configurations. Because time-reversal symmetry remains unbroken at the multi-phase point, we set all visible and hidden biases to zero, leaving the coupling matrix $\mathbf{W}$ as the sole set of trainable parameters. To monitor training progress, we track the iteration-to-iteration change in $\mathbf{W}$, quantified by
[…]
Fig. 3 plots $\varepsilon_{W}$ as a function of the iteration number $k$, providing a convenient diagnostic of convergence behavior. At early stages, $\varepsilon_{W}$ exhibits relatively large fluctuations, reflecting the rapid adjustment of the weights as the model aligns itself with the MC-sampled spin configurations. Within the first few tens of iterations, the amplitude of these fluctuations decreases markedly, indicating that the data-dependent and model-dependent contributions to the gradient are approaching balance. Beyond this transient regime, $\varepsilon_{W}$ decays smoothly toward zero over hundreds of iterations, eventually reaching a small, nearly constant plateau. This saturation signals that the RBM parameters have entered a steady-learning regime in which subsequent updates become minimal, consistent with convergence of the stochastic gradient descent procedure.
To benchmark the performance of the trained RBM, we consider the spin-spin correlation function defined as
$C(r)=\langle\sigma_{i}\sigma_{i+r}\rangle-\langle\sigma_{i}\rangle^{2}$ (12)
where the average $\langle\cdots\rangle$ denotes both a spatial average over spin-pairs separated by $r$ within each configuration and an ensemble average over independently sampled configurations. As shown in Fig. 4, the correlation functions computed from RBM-generated samples match those obtained from MC simulations with high fidelity. This agreement demonstrates that the RBM successfully learns not only the uniform probability measure over the degenerate ground-state manifold but also the nontrivial short-range, oscillatory correlations characteristic of the multiphase point $\kappa=1/2$. In particular, reproducing the oscillatory and exponentially decaying $C(r)$ indicates that the RBM is able to internalize the subtle constraints that define the ground-state manifold—despite the absence of explicit long-range order—highlighting its capability to model constrained, frustration-induced statistical distributions.
IV Kagome Spin Ice
Next, we apply the RBM framework to learn the correlated, liquid-like phases of kagome spin ice. Kagome spin ice [56] can be viewed as an antiferromagnetic Ising model on the kagome lattice—a two-dimensional network of corner-sharing triangles; see Fig. 5. In artificial realizations, elongated single-domain ferromagnetic nanomagnets serve as mesoscopic Ising degrees of freedom and are arranged in planar arrays that interact via long-range dipolar fields [62, 63, 64, 65, 66]. In the kagome variant, these nanowires are patterned along the edges of a honeycomb network, the dual of the kagome geometry [67, 68, 69, 70, 71, 72, 73, 74, 75, 76].
Kagome-ice physics also emerges in several bulk magnets. In canonical pyrochlore spin-ice compounds—Dy2Ti2O7, Ho2Ti2O7—an external field applied along [111] pins one of the four sublattices, producing an effective kagome layer whose remaining spins satisfy kagome-ice rules [77, 78, 79]. More recently, intrinsic kagome-ice behavior has been identified in the tripod-kagome compound Dy3Mg2Sb3O14 [80, 81, 82] and in the quasi-two-dimensional intermetallic magnet HoAgGe [82], both hosting frustrated Ising moments on kagome layers.
In what follows, we focus on the two degenerate ice manifolds characteristic of kagome spin ice. The ice-I phase is defined solely by local “ice rules” that enforce a two-in/one-out or one-in/two-out configuration on every triangle. The ice-II manifold is further constrained by the emergence of a long-range ordering pattern of emergent magnetic charges, leading to a more strongly correlated yet still liquid-like phase.
IV.1 Ice-I phase
Figure 6: Schematic illustration of a RBM for learning the ice phases on the kagome ice. The visible layer represents the Ising spins on the kagome lattice, while the hidden layer encodes latent features associated with the ice-rule constraints and short-range correlations. Trainable parameters include the visible and hidden biases (local fields) and the visible-hidden coupling weights, all of which are optimized using MC-generated spin configurations.
To formulate the kagome spin-ice Hamiltonian, we define unit vectors $\hat{\mathbf{e}}_{i}$ pointing along the local Ising axes from the center of each up-triangle to its corners. Each magnetic moment is written as $\mathbf{S}_{i}=\sigma_{i}m_{0}\hat{\mathbf{e}}_{i}$, with $\sigma_{i}=\pm 1$ denoting its Ising orientation. Because nearest neighbors satisfy $\hat{\mathbf{e}}_{i}\cdot\hat{\mathbf{e}}_{j}=-1/2$, a ferromagnetic exchange $J_{F},\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ ($J_{F}<0$) produces an antiferromagnetic Ising coupling,
$\mathcal{H}_{\mbox{\scriptsize ice-I}}=J\sum_{\langle ij\rangle}\sigma_{i}\sigma_{j}=\frac{v}{2}\sum_{\alpha}Q_{\alpha}^{2}+E_{0},$ (13)
with $J=|J_{F}|/2$. The second form rewrites the Hamiltonian in terms of magnetic charges $Q_{\alpha}$ on triangles. A convenient picture is the dumbbell model [83], where each dipole is replaced by a bar magnet of length $\ell$ whose poles of charge $q_{m}=m_{0}/\ell$ sit at neighboring triangle centers. The net charge on triangle $\alpha$ is
$Q_{\alpha}=\pm q_{m}\sum_{i\in\alpha}\sigma_{i},$ (14)
with the sign distinguishing up- and down-triangles. In these units, $v=J\ell^{2}/m_{0}^{2}$. Minimizing Eq. (13) requires each triangle to carry the smallest possible charge, $Q_{\alpha}=\pm 1$, corresponding to 2-in/1-out or 1-in/2-out spin arrangements, as shown in Fig. 5(a). These are the kagome-ice analog of the Bernal-Fowler ice rules [84]. The correlated states defined by these 2-in/1-out or 1-in/2-out constraints constitute the ice-I manifold. Their exponentially large number yields a finite residual entropy per spin, $S_{0}\approx\frac{1}{3}\ln(9/2)$ [56]. The magnetic-charge representation in Eq. (13) also makes evident that excitations out of this manifold correspond to defect triangles with $Q=+3$ or 3-3, i.e., 3-in or 3-out configurations; see Fig. 5(b). These defects act as magnetic monopoles within the kagome-ice framework and control the low-energy excitations above the ice-rule manifold. A representative spin-disordered ice-I state, including an example of a $Q=\pm 3$ defect pair, is shown in Fig. 5(c).
Here we implement an RBM to learn the ice-I manifold of kagome spin ice; a kagome-specific schematic is shown in Fig. 6. The visible units of the RBM encode the Ising spins on the kagome lattice in a fixed ordering, while effective interactions among these spins are mediated through their couplings to the hidden layer, as reflected in the effective energy expression Eq. (7). As discussed in Sec. II, training adjusts the RBM parameters so that the marginal distribution Eq. (5) provides an accurate approximation to the zero-temperature ground-state manifold of Eq. (13), i.e. $\pi_{\bm{\theta}}(\bm{\sigma})\approx\pi_{\scriptsize\mbox{ice-I}}(\bm{\sigma};\,T=0)$, which is a uniform distribution over all configurations satisfying the local ice rules. Because nonzero bias fields $\bm{b}$ and $\bm{b}^{\prime}$ explicitly break time-reversal symmetry, $\bm{\sigma}\to-\bm{\sigma}$, and the ice-I manifold is itself time-reversal invariant, we set both biases to zero throughout training.
Figure 7: Comparison of the spin-spin correlation function $C(r)$ obtained from MC sampling and from RBM-generated configurations for the $T=0$ ice-I manifold of $\mathcal{H}_{\scriptsize\mbox{ice-I}}$. The inset shows $|C(r)|$ on a semi-logarithmic scale, highlighting the short-range nature of the correlations. The dashed line indicates an exponential fit, $\exp(-r/\ell)$, with correlation length $\ell\approx 0.95$.
The training data for the kagome ice-I phase were generated using standard Markov-chain Monte Carlo (MC) simulations on a $15\times 15$ lattice (with $N=3\times 15^{2}$ spins) at $T=0$. In contrast to the more constrained pyrochlore ice or the charge-ordered ice-II manifold discussed below, the kagome ice-I phase possesses an extensive set of local zero modes: many single-spin flips preserve the 2-in/1-out (or 1-in/2-out) constraint on every triangle. As a result, single-spin Metropolis updates efficiently explore the ice-I manifold without becoming trapped in restricted sectors of configuration space. As shown in Fig. 7, an RBM trained on these MC samples accurately reproduces the spin–spin correlation function, indicating that it has learned the essential statistics of the ice-I phase.
The correlation function itself reflects the characteristic short-range order imposed by the ice rules. The oscillatory behavior with a period of one lattice spacing arises from the underlying nearest-neighbor antiferromagnetic interaction of the Ising spins, while the strong overall exponential envelope, $e^{-r/\ell}$, evidences the geometric frustration that prevents long-range order. This exponential decay is clearly visible in the semi-logarithmic inset of Fig. 7, with a correlation length $\ell\sim\mathcal{O}(1)$ lattice spacing. The agreement between RBM-generated data (red circles) and MC results (blue triangles) across all distances demonstrates that the trained RBM successfully captures both the local ice-rule constraints and the rapidly decaying correlations characteristic of the kagome ice-I manifold.
IV.2 Ice-II phase
In real kagome-ice materials, long-range dipolar interactions play a central role. In artificial kagome ices in particular, the nearest-neighbor coupling $J$ that stabilizes the ice-I phase arises entirely from the dipole-dipole interaction between magnetic nano-islands. While the nearest-neighbor model supports a massively degenerate ice-I manifold, dipolar interactions lift this degeneracy and generate additional correlations. At low temperature they eventually drive true long-range order, but at intermediate temperatures they stabilize an emergent ice-II phase: a partially ordered state in which the residual charges $Q_{\alpha}=\pm 1$ arrange in an alternating pattern on the dual honeycomb lattice.
This charge ordering is most naturally interpreted through the effective Coulomb interaction between triangle charges, $(\mu_{0}q_{m}^{2}/8\pi)\,Q_{\alpha}Q_{\beta}/|\mathbf{r}_{\alpha}-\mathbf{r}_{\beta}|$, which gives the leading-order approximation of the dipolar interaction. To capture the essential physics of the ice-II phase within a minimal theoretical framework, we consider a spin-ice Hamiltonian supplemented by a short-range interaction between neighboring triangle charges:
$\mathcal{H}_{\mbox{\scriptsize ice-II}}=\frac{v}{2}\sum_{\alpha}Q_{\alpha}^{2}+K\sum_{\alpha\beta}Q_{\alpha}Q_{\beta},$ (15)
where Greek indices label the triangular plaquettes of the kagome lattice. The coefficient $v=J\ell^{2}/m_{0}^{2}$ represents the local self-energy of magnetic charges generated by the nearest-neighbor antiferromagnetic exchange $J$. Minimizing this term enforces $Q_{\alpha}=\pm 1$, recovering the 2-in/1-out (or 1-in/2-out) ice rule discussed in Sec. IV.1. The second term, proportional to $K>0$, introduces an antiferromagnetic interaction between charges on adjacent triangles. This repulsive coupling selects a staggered pattern of $Q_{\alpha}=\pm 1$ on the honeycomb dual lattice, thereby breaking the $Z_{2}$ symmetry: all up-triangles acquire the same charge (say $Q=+1$), while all down-triangles carry the opposite charge. This simple extension thus encapsulates the emergence of charge order characteristic of the kagome ice-II phase.
Importantly, the spins themselves remain disordered within the charge-ordered state; see Fig. 5(c) for an example. The residual degeneracy can be quantified by mapping every charge-ordered configuration to a dimer covering of the honeycomb lattice: in each triangle with $Q_{\alpha}=\pm 1$, two majority spins point into (or out of) the triangle, and the minority spin points in the opposite direction. Associating these minority spins with dimers on the dual honeycomb lattice establishes a one-to-one correspondence with perfect dimer coverings. Because the number of honeycomb dimer coverings grows exponentially with system size, the ice-II phase retains a finite residual entropy, $S=0.108$ per spin [85]. The ice-II phase therefore exemplifies ordering driven by emergent degrees of freedom—magnetic charges—rather than by the microscopic spins themselves, and highlights the hierarchical structure of correlations in kagome spin ice.
To train RBMs for the ice-II phase, we performed MC simulations of the Hamiltonian in Eq. (15) at temperatures $T\lesssim K$, where the short-range charge interaction stabilizes staggered charge order. Throughout, the temperature $T$ is measured in units of the nearest-neighbor coupling $J$, and we set the charge–charge repulsion to $K=0.1J$. This order imposes additional constraints beyond the ice rules and greatly restricts the accessible configuration space. In a charge-ordered background, a single-spin flip almost always disrupts the $Q_{\alpha}=\pm 1$ pattern, creating either costly $Q=\pm 3$ monopoles or the less severe but still unfavorable misplaced $Q=\pm 1$ defects. As a result, naive single-spin Metropolis updates become inefficient. To ensure adequate sampling, our MC simulations therefore combine local single-spin flips with loop updates that flip spins along closed paths while preserving the underlying charge pattern.
The emergence of charge order also breaks a $Z_{2}$ symmetry associated with time-reversal. This implies that the effective RBM Hamiltonian in Eq. (7) must explicitly break time-reversal symmetry, highlighting the importance of including local bias fields. Indeed, unlike in the ice-I case, RBMs trained without local fields perform poorly for ice-II configurations. Therefore, in the RBM training discussed below, we treat $\bm{b}$, $\bm{b}^{\prime}$, and $\bm{W}$ as full sets of trainable parameters. These parameters are adjusted through the training process such that the marginal distribution Eq. (5) accurately approximates the Boltzmann distribution of the ice-II Hamiltonian at the given temperature, i.e. $\pi_{\bm{\theta}}(\bm{\sigma})\approx\pi_{\scriptsize\mbox{ice-II}}(\bm{\sigma};\,T)\propto\exp(-\mathcal{H}_{\scriptsize\mbox{ice-II}}/T)$.
Figure 8: Comparison of the spin-spin correlation function $C(r)$ obtained from Markov-chain MC simulations and RBM-generated configurations for the kagome ice-II phase at $T=K=0.1J$. The inset shows the absolute value $|C(r)|$ on a log-log scale, where the dashed line indicates a $1/r^{2}$ power-law decay.
Fig. 8 compares the spin-spin correlation function $C(r)$ of the kagome ice-II phase obtained from Markov-chain MC simulations with that generated by a trained RBM. The presence of the staggered charge order enhances the short-range antiferromagnetic correlations between spins. the RBM-sampled values (red circles) track the Monte Carlo results (blue triangles) semiquantitatively, capturing the oscillatory behavior and the rapid suppression of correlations that follow from the disordered spin background. The overall agreement demonstrates that, once equipped with symmetry-breaking bias fields, the RBM successfully learns both the local and intermediate-range structure of the ice-II phase.
The inset of Fig. 8 presents $|C(r)|$ on a log-log scale, revealing a clear power-law envelope. The dashed line corresponds to a $1/r^{2}$ power-law decay, in agreement with the expected asymptotic form for correlations in the ice-II manifold, where long-range charge order coexists with algebraically decaying spin correlations inherited from an underlying dimer-like description [86]. That the RBM reproduces this power-law behavior—despite the absence of explicit long-range couplings in its architecture—indicates that the hidden layer has effectively learned the emergent constraints imposed by the staggered charge pattern.
Figure 9: Histogram of RBM weights $W_{ij}$ for ice-I (red) and ice-II (blue) phases. The ice-I distribution is broad and symmetric, reflecting the highly fluctuating ice manifold, whereas the ice-II weights are sharply peaked near zero, consistent with the reduced configurational variability in the charge-ordered phase.
The histogram of trained RBM weights provides a useful window into how the model encodes the distinct correlation structures of kagome ice-I and ice-II. As shown in Fig. 9, the weight distribution for the ice-I manifold (red) is broad and nearly symmetric, with substantial support for both positive and negative couplings. This reflects the fluctuating character of ice-I: the two-in/one-out constraint enforces only short-range correlations and leaves a macroscopically degenerate configuration space, which the RBM captures through a diffuse set of visible-hidden couplings.
By contrast, the distribution learned for the ice-II phase (blue) is significantly narrower and more sharply peaked near zero. Long-range charge order reduces the variability of local spin environments, allowing the RBM to encode the data using smaller-magnitude couplings. The pronounced narrowing of the weight histogram thus mirrors the reduced configurational entropy and enhanced regularity of the charge-ordered phase. These weight statistics therefore offer a direct diagnostic of the underlying correlation hierarchy, naturally distinguishing the extensively degenerate ice-I manifold from the partially ordered ice-II phase.
Figure 10: Histograms of the RBM local fields for the visible (light curves) and hidden (dark curves) layers at two temperatures within the ice-II phase: (a) $T=1.0$ and (b) $T=0.1$, in units of the nearest-neighbor AF coupling $J$ with charge–charge repulsion fixed at $K=0.1J$. At both temperatures, the visible-layer fields cluster around a single positive sign, while the hidden-layer fields peak around the opposite sign, consistent with the broken time-reversal symmetry of the charge-ordered ice-II phase.
The distributions of the local bias fields learned by the RBM provide a direct probe of how the model captures the broken symmetries and correlation structure of the ice-II phase. As shown in Fig. 10 for two temperatures within the ice-II regime, the visible-layer fields $\bm{b}$ (light curves) are consistently concentrated on the same sign, whereas the hidden-layer fields $\bm{b}^{\prime}$ form a distribution centered at the opposite sign. This clear sign separation persists at both $T=1.0$ and $T=0.1$, indicating that it is a robust feature of the trained models rather than a temperature-dependent artifact. The magnitude of the fields narrows as temperature decreases, reflecting the reduced configurational fluctuations deep in the charge-ordered phase, but the relative sign structure remains unchanged.
The emergence of uniformly signed visible-layer fields is particularly significant. In the RBM effective energy, these fields act as local “magnetic fields” that bias the visible spins toward one orientation. A uniform sign in the visible-layer fields acts as an effective global bias on the physical spins and directly reflects the broken time-reversal $Z_{2}$ symmetry of the ice-II phase. Once monopole charges order, the local spin environments acquire a preferred orientation, and the RBM encodes this by assigning visible biases of consistent sign. The hidden-layer fields show the complementary structure: they peak around the opposite sign, indicating that the hidden units activate in a way that reinforces the broken-symmetry pattern. This opposing-sign structure in the two layers demonstrates that the RBM internalizes the essential symmetry breaking of ice-II and represents it through a coherent organization of its bias fields.
V Conclusion and Outlook
In this work we have demonstrated that Restricted Boltzmann Machines (RBMs) offer a powerful and flexible generative framework for learning the statistical structure of frustrated magnetic systems. As a controlled benchmark, we first applied the RBM to the one-dimensional ANNNI model at its multiphase point, where the ground state is highly degenerate and characterized by oscillatory, exponentially decaying spin-spin correlations. The trained RBM successfully reproduced these features with quantitative accuracy, showing that even a shallow generative model is capable of capturing nontrivial emergent correlations arising from competing interactions.
Building on this foundation, we employed RBMs to learn the spin configurations of kagome spin ice, whose ice-I manifold exhibits extensive degeneracy and short-range correlations enforced by local 2-in/1-out constraints. Using Monte Carlo samples as training data, the RBM faithfully captured these ice rules and accurately reproduced the spin-spin correlation functions of the ice-I phase. We further examined the more intricate ice-II phase, where long-range charge order emerges from effective Coulomb interactions between magnetic charges. Here, we found that correct modeling requires RBMs with nonzero, uniform-sign bias fields, reflecting the underlying breaking of time-reversal symmetry. Together, these results demonstrate that RBMs are capable of learning both local constraints and emergent correlations in frustrated magnets, even when the underlying manifolds are highly degenerate or partially ordered.
Our findings highlight several promising directions for future work. Extending the present approach to larger system sizes and deeper or more expressive architectures—such as convolutional RBMs and tensor-network inspired deep Boltzmann machines—may further enhance the model’s ability to capture long-range emergent structures in more complex frustrated systems. Finally, the success of RBMs in learning kagome spin-ice manifolds suggests that generative models may serve as useful tools for probing otherwise intractable frustrated systems, including classical spin liquids, Coulomb-phase systems, and systems with gauge constraints. By providing compact probabilistic representations of these highly correlated states, generative models hold the potential to complement conventional simulation methods and offer new perspectives on emergent phenomena in frustrated magnetism.
Acknowledgements.
This work was supported by the US Department of Energy Basic Energy Sciences under Award No. DE-SC0020330. The authors also acknowledge the support of Research Computing at the University of Virginia.